User Tools

Site Tools


This wiki is not maintained! Do not use this when setting up AuScope experiments!

Chair: Mas

Attended: Jamie, Lucia, Mas, Simin, Katie, Tiege, Lucas, Patrick, Ross, Artifur

Apologies: Johnny

Absent: Arwin

Handover Issues


Katherine 12m

•Patrick need to swapped out the module due to bad disk, received an alarm error (m5-104 mk5cn: time-out, connection closed), tried to restart the mk5 in between scans but didn’t fix the error. The error is due to a faulty disk in the module (shows up in get_stats procedure), which causing writes to disk to be slower than usual. Have asked Mick to swap to a good module in the morning.

•Patrick received the following alarms (ALARM: error st -998 reading systemclock1, ALARM: error st -999 tcp/ip connection error, and WARNING: error st -10 ramp epoch time not -1 when preparing to load). Has checked antenna and it is in “operate” mode.

•Received WARNING: error st -5 error return from antenna and see mbus error. Have to halt and stow the antenna twice to fix it. In the morning a new module was swapped.

–Mk5 was responding. Mick swapped the module. Can continue to observe–


Yarragadee 12m

•Katie received “ALARM: error m5-104 mk5cn: time-out, connection close” continuously. Had called Mick and he said only 1 light was on for the drives. Looks like mk5 crashes and had to reboot it.

•Jamie restarted pc field system, and had to restart e-remote control as it lost connection (Need to reboot)


Hobart 26m

•Lucas noticed that the disk position has been drifting very quickly, 360 GBs in 4.5 hours, aka 80GB/hour.

–Being scheduled by HartRao using their own software. Summary file is not trustworthy–

•Simin received regular “ALARM: error f1 -1 previous source in this schedule not reached before new source was commanded” and several missed scans

– windstow the telescope and flood occurred–

Katherine 12m

•Lucas noticed that the disk position has been drifting, 25 GBs in 4.5 hours, aka5.5GB/hour (Lucas).

•Simin received several missed scans

Yarragadee 12m

•Lucas received constant onsource slewing errors, and it seems like Yarragadee is late to source quite often. The disk position has been drifting very quickly, 100 GBs in 4.5 hours, aka 22GB/hour.


Katherine 12m

•Warren noticed large delay and reconfigure DBBC

•Arwin noticed delay was drifting again and power-cycled DBBC


Katherine 12m

•Jamie halted the schedule to restart DBBC. Clkoff had been slipping and passed the threshold. Several data were affected by delay jumps.

–Brett and warren working on cable wrap–

–Warren and Brett are working for Katherine issue, so need people to go to Mt Pleasant during daytime–

–Katherine is out, only Yarragadee left–

–Phone in auscope room is not working, use Skype, can dial out on line, no one can call in–

–Next Tuesday session start early–

–Please switch the alarm when you swap with observer who observe at Mt Pleasant (close the log monitor)–

–Jamie and Lucia are not around in June. Only Warren is left as oncall person–

–Volunteers for the Wiki???–

/home/www/auscope/opswiki/data/pages/operations/agenda2018_05_17.txt · Last modified: 2018/05/17 03:22 by Noor Md Said